x2(x squared) + 16=8x
I changed your problem. See the solution section below for an explanation.

Answer by mathsmiles(68)

x^2 + 16 = 8x
First, we need to move everything to one side of the equation and set it equal to zero.
x^2 + 16 = 8x Subtract 8x from both sides:
x^2 - 8x + 16 = 0 Now factor this. We're looking for two factors of 16 whose difference is 8. Hmm... Are you sure you didn't copy the problem down wrong and it's
x^2 + 16 = -8x (minus 8x)? I bet it is, so I'm going to solve that instead:
x^2 + 16 = -8x Add 8x to both sides
x^2 + 8x + 16 = 0
(x + ...)(x + ...) = 0 Now find two factors of 16 whose sum is 8. That would be 4x4=16 since 4+4= 8
(x + 4) (x + 4) = 0
At least one of these factors needs to be zero for this to be true:
(x+4) = 0 | (x+4) = 0
Subtract 4 from both sides:
x = -4
Checking:
x^2 + 16 = -8x
(-4)^2 + 16 = -8(-4)
16 + 16 = 32
32 = 32 Check!